The E can be electrolyzed  after melting  to yield 4.0 grams of F and 1.6 grams of G,   neither  of which can be further decomposed by ordinary chemical means.  E combines  with water to give 1.3 grams of J per gram of E and gas L. The gas L has exactly the same    properties as the product obtained when carbon is burned in an excess of oxygen.

1
Expert's answer
2021-09-19T00:03:26-0400

4/1.6 = 2.5

1.6/1.3 = 1.23

Gas L = 2.5/1.23= 2.03/2,= 1g